### v1

# IceCube Collaboration

Abstract
We report a measurement of the inelasticity distribution in the scattering of neutrinos of energy $80-560~{}\mathrm{GeV}$ off nucleons, which is sensitive to the inclusive differential cross section. This analysis is based on a sample of atmospheric muon neutrinos detected in the IceCube sub-array DeepCore during 2012-2021, and is the first such measurement in this energy range. Our measurement extends to energies where accelerator data is not available, hence we compare our results to predictions from perturbative QCD calculations, finding good agreement.
footnotetext: email: analysis@icecube.wisc.edu (corresponding author)
I Introduction
[…]
Another source of information on neutrino interactions are neutrino telescopes which detect naturally created neutrinos. These experiments access much higher energies, but are limited by the uncertainties on the incoming flux and the accuracy with which an interaction can be reconstructed. Despite these limitations, these experiments can contribute in a complimentary way to the understanding of neutrino cross sections. IceCube, which detects both atmospheric and astrophysical neutrinos, has previously reported measurements of the total neutrino cross section up to a PeV [5, 6], as well as a measurement of the inelasticity distribution at neutrino energies of $1-20$ TeV [7]. Such energies are well above these accessible at terrestrial accelerators.
Here, we present the results from a measurement of the flux-averaged inelasticity distribution in neutrino-nucleon interactions from 80 GeV to 560 GeV using data from the IceCube DeepCore array. At these energies neutrinos interact mainly through deep-inelastic scattering (DIS), which involves scattering off a quark or gluon in the nucleon, resulting in a lepton and a hadronic shower in the final state. The inelasticity $y$ is a measure of the energy distribution of the products of the interaction, defined as [8]
$y=1-E_{\mathrm{lepton}}/E_{\nu},$ (1)
where $E_{\nu}$ is the energy of the incoming neutrino and $E_{\mathrm{lepton}}$ is the energy of the outgoing lepton. In this study, we follow Ref. [7] and fit an empirical model of the inelasticity distribution shape to atmospheric neutrino data at different values of $E_{\nu}$, rather than doing a differential cross section measurement.
A measurement of the inelasticity of neutrino-nucleon interactions simultaneously probes the fundamental properties of the weak force and our understanding of the structure of the nucleon in QCD. At energies below 1 PeV neutrinos and antineutrinos have different inelasticity distributions as the bulk of their interactions happen on valence quarks. The difference disappears at high energies as neutrinos and antineutrinos are able to interact with equal numbers of sea quarks and anti-quarks and with gluons [8]. The measurement presented here uses the atmospheric neutrino flux, which contains both neutrinos and antineutrinos. We measure a convolution of the neutrino and antineutrino inelasticity distributions with the neutrino to antineutrino ratio. Such a measurement can validate current cross section and flux models, while any deviation in the measured inelasticity distribution could indicate a mismodeling either in cross section or in atmospheric flux predictions. We also invert our results to measure the neutrino fraction in our sample, assuming a specific cross section calculation.
At $E_{\nu}=$ 80-560 GeV, neutrino-nucleon interactions are expected to be dominated by DIS, however, resonance (RES) and quasi-elastic (QE) scattering events provide a non-negligible contribution at low inelasticity values. In this study we measure the inclusive inelasticity distribution. This is because in DeepCore we cannot reliably differentiate between QE, RES and DIS events due to the large inter-module spacing compared to the characteristic size of the hadronic cascades.
We begin by introducing the IceCube Neutrino Observatory (Sec. II), followed by a description of the data sample used (Sec. III). The analysis is presented in Sec. IV, and results in Sec. V. In Sec. VI we discuss the implications of our results. The limitations of the parametrization used in our analysis are described in detail in Appendix A. In Appendix B we show a detailed comparison of the impact of measured and nuisance parameters on our data sample.
II The IceCube neutrino detector
The IceCube Neutrino Observatory is an ice-Cherenkov detector consisting of 5,160 Digital Optical Modules (DOMs) deployed deep under the surface at the geographic South Pole. The DOMs instrument a volume of approximately 1 km3, starting at a depth of 1,450 m below the surface and ending at 2,450 m. The bulk of the array is optimized for the observation of neutrinos with energies of 1 TeV and above, and has DOMs spaced by 17 m in the vertical direction, and 125 m in the horizontal direction [9]. A low-energy extension, known as DeepCore, is located at the bottom, central region of the detector, where higher quantum efficiency PMTs are deployed closer together, with a vertical separation of 7 m and a horizontal separation between 40 m and 70 m [10]. DeepCore can observe neutrino events down to $\sim$10 GeV [11].
The sensitive element in the DOMs is a 10-inch photomultiplier tube (PMT) facing downward away from the ice surface [12]. The PMT is enclosed in a glass sphere together with the electronics for power and signal digitization [9]. The waveforms produced by the PMTs in each DOM are recorded by analog-digital converters and the digitized data is passed to a pulse-finding algorithm, that deconvolves it into pulses using the typical response of the PMT to single photons [13, 14]. The interpretation of the properties of the events, such as the event selection and reconstruction described below, is carried out based on the information contained in these pulses.
In this analysis we use DeepCore to study the inelasticity of events with primary neutrino energies in the range from $80$ to $560$ GeV. The hadronic cascade in such events are still mostly contained inside the more densely instrumented DeepCore volume, which gives us a more precise reconstruction compared to the main IceCube array. The energy of the muons exiting the DeepCore, but still contained inside the IceCube, can be estimated from their track length.
III Data selection
The type of events that are most useful for this analysis are those where the outgoing lepton can be distinguished from hadronic interaction products, and the inelasticity can therefore be estimated. This limits us to $\nu_{\mu}$-nucleon charged-current (CC) interactions, where the products are a hadronic shower starting at the interaction point and developing over a distance of meters, and a muon that travels roughly 4.5 m per GeV of its kinetic energy.
The data sample used for this analysis comes from the common DeepCore event selection described in Ref. [11], specifically at Level 5. The common selection focuses on rejecting events produced by atmospheric muons as well as detector noise, while keeping events where the neutrino interacted inside the DeepCore region. The sample after the common selection has approximately a $2:1$ ratio of neutrinos to non-neutrino backgrounds, so further levels of selection described in this section, are necessary for refining a high purity $\nu_{\mu}$ CC-only sample for this study.
III.1 Event reconstruction and classification
Selected events are reconstructed using RETRO, an IceCube low-energy reconstruction method described in Ref. [15]. We reconstruct this sample under the hypothesis that the interaction produces a hadronic shower and a muon, each with a distinct energy, but traveling in the same direction. This hypothesis is reasonable for $\nu_{\mu}$ CC interactions with $E_{\nu}>10$ GeV, the events that are most useful for this study. The inelasticity can therefore be estimated as
$y_{\mathrm{reco}}=1-E_{l,\mathrm{reco}}/(E_{l,\mathrm{reco}}+E_{h,\mathrm{reco}}),$ (2)
where $reco$ stands for reconstructed, $E_{l}$ is the energy of the lepton, and $E_{h}$ is the total energy of the hadronic shower.
Events in 9.28 years % of sample
DIS $\nu_{\mu}$ CC 4756 68.3 %
DIS $\bar{\nu}_{\mu}$ CC 1965 28.0 %
Non-DIS $\nu_{\mu}+\bar{\nu}_{\mu}$ CC 188 2.7 %
$\nu$ NC 27 0.4 %
$\nu_{e}+\bar{\nu}_{e}$ 20 0.3 %
Atm. $\mu$ 11 0.2 %
$\nu_{\tau}+\bar{\nu}_{\tau}$ 10 0.1 %
Noise 0 0.0 %
Table 1: Sample composition at the final level of selection, as described in Sec. III, with projected event counts for 9.28 years of data, calculated using the baseline Monte-Carlo set.
After the reconstruction is performed, two boosted decision tree (BDT) classifiers are applied to the events. Both classifiers were trained on simulation and rely on reconstructed variables [16]. The first is Muon Classifier, which is trained to identify remaining atmospheric muons in the sample. It uses inputs of five variables, namely the Level 4 Muon Classifier score [11], vertical and radial position of the reconstructed interaction vertex, vertical location of the deepest pulse recorded in the ice which matches a hypothesis of muon traveling in a “corridor” between strings, and the difference between reconstructed directions using RETRO and SANTA reconstruction algorithms [15]. A comparison of Muon Classifier score for data and simulation is shown on Fig. 1.
The second BDT classifier determines the Particle Identification Score (PID), which divides events into track-like and cascade-like. The classifier is trained to identify $\nu_{\mu}$ CC interactions as track-like events, while $\nu_{e}$ CC interactions represent cascade-like events. The BDT uses input of five variables coming from the reconstruction algorithm: track length of the outgoing muon, cascade energy, zenith angle, zenith angle uncertainty, and the likelihood difference between the best fit hypothesis with arbitrarily long track and with track length fixed at zero. A comparison of PID Classifier score for data and simulation is shown on Fig. 2, which also indicates the cut on PID score $>$ 0.7 described in Sec. III.3.
Figure 1: Muon Classifier score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the cut on Muon Classifier score, with values to the left of the line removed from the final sample.
Figure 2: PID score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the cut on PID score, with values to the left of the line removed from the final sample.
Figure 3: Level 4 Noise Classifier score comparison for data and for predicted sample composition at the reconstruction level. The Monte-Carlo normalization is scaled to match total event counts to data. The grey dashed line indicates the position of the additional cut on the classifier score, with values to the left of the line removed from the final sample.
Figure 4: Reconstruction resolutions for neutrino energy (left) and inelasticity (right) at the final level of event selection.
In order to further reduce non-neutrino background, the Muon Classifier score is required to be $>0.4$ for events to pass to the next step of selection. An additional cut of $>0.95$ is applied on the Level 4 Noise Classifier [11] to reduce remaining background. The performance of the Level 4 Noise Classifier score for our sample at the reconstruction level is shown in Figure 3.
III.2 Rejection of non-neutrino backgrounds
Another class of variables used to remove noise and low quality events are calculated based on estimated photon speeds, under the assumption that each detected photon has started at the event interaction vertex and and then travelled without scattering to the corresponding DOM. Photon speeds estimated in this way will not necessarily accurately represent the real event, however the variables calculated using these are still useful, as they provide information on the event quality. In our event selection, we require the median estimated speed of all detected photons in an event to be $>0.4c$, where $c$ is the speed of light in vacuum. We also require that in each event the fraction of detected photons with estimated speeds below zero or above $c$ is $<0.4$, because such pulses typically result from the PMT noise. A photon speed below zero indicates that the pulse happened before the reconstructed interaction time, while photon speed above $c$ means that the pulse happened either too far from the reconstructed vertex or too early to be explained by a photon travelling from the vertex at the speed $c$. While a photon travelling in ice will have a speed of $c/n$, where $n$ is the refractive index of ice, the choice was made nevertheless to make a cut at the speed of $c$, to avoid penalizing events where a muon travels at speed higher then $c/n$ and than produces a Cherenkov photon near the DOM.
Restrictions are also applied to reject neutrino events with likely coincident atmospheric muons entering the detector. This is achieved by introducing cuts on events with pulses in the veto region. This region includes the top and the outermost optical modules of IceCube array, as these are more likely to detect light from atmospheric muons entering the detector. We limit the allowed number of pulses in the outermost layer of IceCube to be below 8. At the same time we also require that there is no downward travelling trend in the pulse distribution in the top 15 layers of DOMs in IceCube, which corresponds to a vertical range of roughly 255 m.
We introduce vertex containment cuts on reconstructed vertex position to remove events with vertex located outside the DeepCore. We require that the vertical position of the vertex be in the range -500 m $<z_{\mathrm{reco}}<$200 m from the vertical center of IceCube and that horizontal radial distance from the central string be $\rho_{\mathrm{reco}}<$ 300 m.
III.3 Rejection of non-$\nu_{\mu}$ CC backgrounds
The remaining steps focus on selecting $\nu_{\mu}$ CC events with high quality reconstruction. The primary event types targeted for the removal are very dim atmospheric muons, which leave few or no pulses on the veto region, and were therefore missed by previous selection steps, and non-$\nu_{\mu}$ CC neutrino interactions.
Dim muons are rejected by estimating how well the track hypothesis used fits the data. To do this, we require that at least 10 pulses should be observed within a time window of [-10,+900] ns from their expected arrival time assuming Cherenkov emission without scattering. This cut removes an estimated 95% of the muons, while reducing the signal by 20%.
Other neutrino interactions are rejected by applying a strict cut on the PID score, which is required to be higher that $0.7$ for the events to pass the selection. Events with $y_{\mathrm{reco}}<0.001$ or $E_{h,\mathrm{reco}}=0$ GeV are removed, as they were found to contain a large fraction of events with the cascade located either in the less densely instrumented region outside the DeepCore or fully below the detector. Finally, we also introduce a cut on the reconstructed zenith direction $\cos\theta_{z,\mathrm{reco}}<-0.2$ to keep only up-going events, which are likely to illuminate the PMT inside a few DOMs without the need for the light to scatter. Upgoing events are also less likely to contain misidentified atmospheric muons due to the shielding by the earth.
III.4 Final level sample
The composition of the final sample is given in Table 1. Our selection results in a signal purity of 99%. The average ratio of neutrinos to antineutrinos is 2.4, which reflects the properties of the atmospheric neutrino flux, the neutrino cross section, as well as event selection efficiency dependence on true inelasticity. The data does not provide sufficient information to distinguish between neutrinos and antineutrinos, so the analysis is done for the sum of both.
The reconstruction resolutions for neutrino energy and inelasticity in the final simulated sample are presented in Fig. 4. The energy estimator shows a clear correlation with neutrino energy up to 500 GeV. Above that, the resolution degrades, because the typical size of the hadronic cascade approaches the size of DeepCore and a larger portion of the muon track could be located outside DeepCore. The inelasticity estimator also displays a correlation across all values, with deviations observed towards the edges, as is expected from one-sided errors in those regions. These two reconstructed quantities are used to extract the inelasticity measurement, as described in the next section.
IV Analysis method
IV.1 General strategy
Figure 5: Predicted energy distributions for each reconstructed energy bin with 9.28 years of data using the GENIE cross section model [17] and HKKMS’15 atmospheric neutrino flux [18]. The shaded bands indicate the central 68% quantile for the energy bin of a corresponding color and the dashed lines indicate the median true energy of the bin.
Figure 6: Expected event count distribution in the analysis template for 9.28 years of data, assuming the GENIE cross section model and HKKMS’15 flux.
Figure 7: Fit of the parametrization (3) to inclusive inelasticity distributions for generator level GENIE Monte-Carlo. Fits were done separately for neutrino and antineutrino inclusive inelasticity distributions in three energy bins, which were weighted to reflect energy distributions in energy bins at analysis level. Fit results for parameters $\langle y\rangle$ and $\log_{10}\lambda$ are shown in the bottom left of each plot.
Figure 8: Data/Monte-Carlo agreement post-fit for the reconstructed inelasticity distribution in each of three $E_{\mathrm{reco}}$ bins.
This analysis is performed by passing data and simulation through the same selection and reconstruction steps, and then comparing their distributions in the space of observables chosen. The shape of the true inelasticity distribution of the simulation is parametrized, so it can be modified, along with nuisance parameters that describe the impact of different sources of uncertainties, in order to find the values that result in the best fit to the data. In this study, we compare a two-dimensional histogram in $E_{\mathrm{reco}}$ and $y_{\mathrm{reco}}$ with 3 bins in reconstructed energy between 100 GeV and 1 TeV and 11 equally sized bins in reconstructed inelasticity, between zero and one.
Figure 5 shows the true energy distribution for each one of the three bins in reconstructed energy, assuming the GENIE 2.12.8 cross section model [17] and HKKMS’15 atmospheric neutrino flux model by Honda et al. [18]. The shaded bands indicate the central 68% quantile, which is used to report the range in which the results are valid, while dashed lines indicate the median energy of the bin. Figure 6 shows the expectation from simulation for event counts in the analysis binning.
IV.2 Parametrization of inelasticity
To parameterize the shape of the inelasticity distribution, our study takes advantage of the expected dependence of the neutrino-nucleon interaction cross section on inelasticity, as described in Sec. I. The functional form, used in previous IceCube studies [7, 19, 20], is the probability distribution of inelasticity and depends on two parameters $\epsilon$ and $\lambda$,
$\frac{dp}{dy}(\epsilon,\lambda)=\frac{1}{\sigma}\frac{d\sigma}{dy}(\epsilon,\lambda)=N\cdot(1+\epsilon(1-y)^{2})\cdot y^{\lambda-1},$ (3)
where $\sigma$ is total cross section for a given energy, $\frac{d\sigma}{dy}$ is the differential cross section as function of inelasticity for a given energy and $N$ is a normalization factor. In order to treat the distribution as a probability density function it is normalized to one, so
$N=\frac{\lambda(\lambda+1)(\lambda+2)}{2\epsilon+(\lambda+1)(\lambda+2)}.$ (4)
Moreover, the parametrization can be transformed from $f(\epsilon,\lambda)$ into $f(\langle y\rangle,\lambda)$ via
$\langle y\rangle=\frac{\lambda(2\epsilon+(\lambda+2)(\lambda+3))}{(\lambda+3)(2\epsilon+(\lambda+1)(\lambda+2))},$ (5)
where $\langle y\rangle$ is the mean of the inelasticity distribution. Since this quantity has a simple physical interpretation, the analysis is done in terms of $\langle y\rangle$ and $\lambda$.
It is important to emphasize that this parametrization is an effective description of the range of “reasonable” inelasticity distribution shapes, however it has a number of limitations described in Appendix A.
The parametrization can be applied to both neutrino and antineutrino inelasticity distributions separately. Figure 7 shows Eq. 3 fitted to muon neutrino and antineutrino expectation from GENIE for three energy bins. However, due to our inability to differentiate between neutrino and antineutrino events, we use the following parametrization to describe the flux-averaged inelasticity distribution at energy $E_{\nu}$,
$\begin{split}\frac{dp}{dy}_{\mathrm{fl.~{}av.}}(E_{\nu})=&~{}\widetilde{\Phi}_{\nu}(\Phi_{\nu},\sigma_{\nu},\mathrm{sel.};E_{\nu})\cdot\frac{dp}{dy}_{\nu}(E_{\nu})\\ &+\widetilde{\Phi}_{\bar{\nu}}(\Phi_{\bar{\nu}},\sigma_{\bar{\nu}},\mathrm{sel.};E_{\nu})\cdot\frac{dp}{dy}_{\bar{\nu}}(E_{\nu}),\end{split}$ (6)
where $\widetilde{\Phi}_{\nu(\bar{\nu})}$ is an energy-dependent fraction of neutrinos (antineutrinos) in our sample, which depends on atmospheric neutrino (antineutrino) flux $\Phi_{\nu(\bar{\nu})}$, the total neutrino(antineutrino)-nucleon cross section $\sigma_{\nu(\bar{\nu})}$, as well as event selection, which is indicated as $\mathrm{sel.}$ in the equation. $\frac{dp}{dy}_{\nu(\bar{\nu})}$ denotes individual neutrino (antineutrino) inelasticity distributions. It is important to highlight that in Eq. 6 $E_{\nu}$ refers to both neutrino and antineutrino energy.
IV.3 Fit of inelasticity distribution to data
The analysis is performed using a forward folding method [11, 21] where the final level simulation is weighted at each combination of parameters tested, and is subsequently used to populate the histogram that is compared to data. We employ an optimization algorithm to select the values of the parameters that will be tested and minimize the function
$-\log\mathcal{L}=\sum_{i=1}^{N_{bins}}(k_{i}\log\mu_{i}-\mu_{i})-\sum_{j=0}^{M_{priors}}\frac{(p_{j}-\hat{p_{j}})^{2}}{2\sigma^{2}_{j}}$ (7)
where the sum runs over all the bins of the histogram $N_{bins}$, $\mu_{i}$ stands for the number of expected events in the $i$-th bin, which depends on the combination of parameters chosen, and $k_{i}$ is the number of observed events. The second term, where the sum is over the number of parameters with Gaussian (half-Gaussian) priors $M_{\mathrm{priors}}$, introduces prior knowledge on some of the fit parameters $p_{j}$ ($j=0,...,M_{\mathrm{priors}}$), penalizing deviations from their best estimates $\hat{p_{j}}$ in terms of their variance $\sigma^{2}_{j}$.
The neutrino simulation was produced using the GENIE Monte-Carlo generator [17] and weighted according to the expected atmospheric neutrino flux. Additional weighting factors were introduced for neutrinos and antineutrinos to produce flat $\frac{dp}{dy}$ distributions, which could then be modified using Eq. 3 for all events in the same $E_{\mathrm{reco}}$ bin. Reweighting all events to the same parametrized distribution means that we will be measuring flux-averaged $\frac{dp}{dy}$ as shown in Eq. 6.
We use the HKKMS’15 atmospheric neutrino flux model [18], and include the effects of neutrino oscillations using the two-flavor approximation ($\nu_{\mu}\rightarrow\nu_{\tau}$) and the parameters reported in Ref. [22], which is sufficient to describe the data since the low energy threshold of this study is $E_{\mathrm{reco}}$=100 GeV. Atmospheric muons were produced using MuonGun, weighted to follow the expected flux from Refs. [23, 24].
Group Parameter Best fit value Nominal value Prior
Physics parameters $\langle y\rangle$ bin 1 $0.45\pm 0.02\pm 0.005$ 0.5 uniform
$\langle y\rangle$ bin 2 $0.46\pm 0.02\pm 0.004$ 0.5 uniform
$\langle y\rangle$ bin 3 $0.45\pm 0.02\pm 0.006$ 0.5 uniform
$\log_{10}\lambda$ bin 1 $0.02^{+0.09}_{-0.12}$ 0.0 half-Gaussian (-0.4)
$\log_{10}\lambda$ bin 2 $0.10^{+0.06}_{-0.05}$ 0.0 half-Gaussian (-0.4)
$\log_{10}\lambda$ bin 3 $0.10^{+0.05}_{-0.04}$ 0.0 half-Gaussian (-0.4)
Flux systematics $\Delta\gamma_{\nu}$ 0.11 0.00 Gaussian ($\pm$ 0.1)
Detector systematics DOM efficiency 1.07 1.00 Gaussian ($\pm$ 0.1)
Ice absorption 1.00 1.00 uniform
Ice scattering 1.01 1.00 uniform
Rel. eff. $p_{0}$0.5991 -0.2674 Gaussian ($\pm$ 0.6)
Rel. eff. $p_{1}$0.02251 -0.04206 Gaussian ($\pm$ 0.12)
Normalization $A_{\mathrm{eff}}$ scale 0.78 1.00 uniform
Table 2: Summary of the best fit values and settings for both the physics and nuisance parameters used in the analysis. All physics parameters are presented with $1\sigma$ uncertainties calculated using Wilks’ theorem. The $\langle y\rangle$ parameters for each energy bin include an additional uncertainty due to limited MC statistics, as described in Sec. IV.4.
The analysis was implemented in the PISA [25] framework. The optimization was done using the differential evolution global minimization method [26] in order to be able to explore the parameter space thoroughly and also to introduce parameter-dependent boundaries, which are needed as described in Sec. IV.2 and Appendix A.
IV.4 Systematic uncertainties impacting the study
The impacts of all the sources of systematic uncertainties that affect this analysis were introduced as correction factors that modify the weight of simulated events. By changing the weights, these factors modify the expectation of all the bins simultaneously, producing a distinct signature in the $E_{\mathrm{reco}}$ vs $y_{\mathrm{reco}}$ space. These signatures differ from the ones produced by changes in the inelasticity distribution, and therefore allow us to perform the measurement accurately, notwithstanding the uncertainties.
We considered all the sources of uncertainties studied in previous DeepCore analyses, using Ref. [11] as our starting point. We studied their impact on the inelasticity parameters by estimating fit mismodeling and impact on the sensitivity using simulations and defined a subset of systematic uncertainties to be included in the fit to the data. The sources of uncertainties that remain are our limited understanding of the scattering and absorption lengths in the Antarctic ice, the effects of the refrozen ice columns on the angular acceptance of the DOMs, the absolute light collection efficiency of the modules, and the normalization and possible corrections to the energy dependence of the atmospheric neutrino flux. While a different set of inelasticity distribution parameters is fitted for each of the three $E_{\mathrm{reco}}$ bins, the nuisance parameters are common for all bins.
The impact of ice and detector properties are obtained by simulating MC sets with different input values, and parameterizing the changes they produce in the bin counts of the analysis histogram. The flux normalization is introduced as a global scaling factor, while the energy dependence is modeled as a power-law modification with a spectral index $E^{\Delta\gamma}$. The impact of the refrozen ice columns is encoded in two parameters that change the angular acceptance of module. The details of the implementation of these parameters can be found in Ref. [11].
Table 2 presents the parameters fitted in the study, including their nominal expectation and the prior knowledge, when applicable. The normalization is given an uniform prior, while the other systematic parameters that encode sources of uncertainty get a weak Gaussian prior. Additional information on the choice of priors is given in Ref. [11].
The limited Monte-Carlo statistics introduce fluctuations in the selection efficiency for neutrinos and antineutrinos, which can potentially lead to a small bias in the fitted mean inelasticity. To account for this, we include an additional error term when reporting the $\langle y\rangle$ results for a given energy bin, such that
$\langle y\rangle=\langle y\rangle_{\mathrm{best~{}fit}}\pm\sigma_{\mathrm{Wilks^{\prime}~{}th.}}\pm\Delta_{\mathrm{MC~{}stat.}}$ (8)
where $\sigma_{\mathrm{Wilks^{\prime}~{}th.}}$ is the $1\sigma$ interval obtained from the likelihood profile using the Wilks’ theorem, and $\Delta_{\mathrm{MC~{}stat.}}$ is the bias in $\langle y\rangle$ obtained using results of an inject-recover test to a sample where original GENIE inelasticity distribution was not modified by reweighting. The sample was fitted using the parametrization in Eq. 3 and the best fit $\langle y\rangle$ value for each bin was compared to the expectation for the same energy bin, calculated using generator-level simulation, with the difference being assigned to the $\Delta_{\mathrm{MC~{}stat.}}$ term.
V Results
We applied the analysis to data from the 2011-2021 seasons, which correspond to 9.28 years of detector livetime. The post-fit Monte-Carlo is in good agreement with the data, with a $p$-value of 9.5%. Figure 8 shows this for $y_{\mathrm{reco}}$ in each of the three $E_{\mathrm{reco}}$ bins. The mean inelasticity is compatible with a constant value $\langle y\rangle$ = 0.45 across all energies, with a $\lambda$ parameter that differs for the lowest energies. The values for all the fit parameters are shown in Table 2. The top panel of Figure 9 shows two-dimensional confidence regions for $\langle y\rangle$ and $\log_{10}\lambda$ in each $E_{\mathrm{reco}}$ bin. The bottom panel on Figure 9 illustrates $\frac{dp}{dy}$ distributions corresponding to the best fit, as well as sampled points within each confidence region. We also show the expectation for our sample using both GENIE [17] and CSMS [27] cross section models, with the HKKMS’15 atmospheric neutrino flux model [18]. The inelasticity distributions show slight differences in shape, with our fit preferring a relatively lower contribution at low $y$, and a relatively higher contribution at high $y$.
Figure 9: Top: Likelihood space of the physical parameters $(\langle y\rangle,\log_{10}\lambda)$ with contours that denote the uncertainty at 1, 2 or 3 $\sigma$. Bottom: Inelasticity distribution shapes as function of inelasticity that results from the fit, also showing the curves that correspond to the points within each confidence region displayed on the top panel in corresponding colors. The grey lines in the background show a representative sample of the values allowed during fit that are beyond 3 $\sigma$. The figure also includes predictions that result from the combination of GENIE (H2O) or CSMS (isoscalar) cross section and the HKKMS’15 atmospheric neutrino flux calculation.
[…]
Figure 12 shows the mean inelasticity compared with results obtained using higher energy interactions in IceCube [7]. Only a subset of the models tested are shown. It important to note that the mean inelasticity prediction is dependent on the event selection, so the results are not directly comparable in detail, but the general trend is consistent.
[…]
Figure 11: Comparison of measured mean inelasticity to predictions calculated using GENIE and CSMS cross sections for the case of no charm production as well as for the nominal case; the predictions are calculated for both HKKMS and daemonflux atmospheric neutrino flux models.
Figure 12: Measured mean inelasticity $\langle y\rangle$ as a function of neutrino energy is compared to the GENIE 2.12.8 $+$ HKKMS 2015 prediction for our event sample. The IceCube 1 TeV-1 PeV inelasticity result [7] and the CSMS [27] $+$ HKKMS 2015 prediction for the corresponding sample are also shown. The highlighted energy range below 245 GeV is where measurements of differential cross section have been made by NuTeV/CCFR [32] and CDHSW [33].
The mean inelasticity can also be interpreted as a measurement of the ratio of neutrinos to antineutrinos in the atmospheric neutrino flux, if we assume a specific cross section. This comparison is shown in Fig. 13, together with the neutrino-antineutrino fraction expected from the same combinations of models discussed before. The spread of the predictions is significantly smaller than the error bars of our measurement.
Figure 13: Neutrino fraction in the atmospheric neutrino flux calculated for our sample assuming different cross section models and compared to theoretical predictions.
VI Discussion
In this study we used the observed energy in track-like and cascade-like byproducts of neutrino candidate events in IceCube to measure the inelasticity of $\nu-$nucleon interactions. We focus on a highly pure sample of $\nu_{\mu}$-CC events between 80 GeV and 560 GeV that interact inside the DeepCore volume. This is the first time that such details of the differential neutrino cross section have been measured at energies between 245 GeV and 1 TeV.
The cross section as a function of inelasticity was parameterized as before [7], and we measured the the two parameters that define its shape $\langle y\rangle$ and $\lambda$. We then compared the mean inelasticity $\langle y\rangle$ obtained with the expectation from three cross section models and two atmospheric neutrino fluxes.
[…]
Acknowledgements.
We thank Yuri Onishchuk for contributions to the early phase of this work. The authors gratefully acknowledge the support from the following agencies and institutions: USA – U.S. National Science Foundation-Office of Polar Programs, U.S. National Science Foundation-Physics Division, U.S. National Science Foundation-EPSCoR, U.S. National Science Foundation-Office of Advanced Cyberinfrastructure, Wisconsin Alumni Research Foundation, Center for High Throughput Computing (CHTC) at the University of Wisconsin–Madison, Open Science Grid (OSG), Partnership to Advance Throughput Computing (PATh), Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS), Frontera and Ranch computing project at the Texas Advanced Computing Center, U.S. Department of Energy-National Energy Research Scientific Computing Center, Particle astrophysics research computing center at the University of Maryland, Institute for Cyber-Enabled Research at Michigan State University, Astroparticle physics computational facility at Marquette University, NVIDIA Corporation, and Google Cloud Platform; Belgium – Funds for Scientific Research (FRS-FNRS and FWO), FWO Odysseus and Big Science programmes, and Belgian Federal Science Policy Office (Belspo); Germany – Bundesministerium für Bildung und Forschung (BMBF), Deutsche Forschungsgemeinschaft (DFG), Helmholtz Alliance for Astroparticle Physics (HAP), Initiative and Networking Fund of the Helmholtz Association, Deutsches Elektronen Synchrotron (DESY), and High Performance Computing cluster of the RWTH Aachen; Sweden – Swedish Research Council, Swedish Polar Research Secretariat, Swedish National Infrastructure for Computing (SNIC), and Knut and Alice Wallenberg Foundation; European Union – EGI Advanced Computing for research; Australia – Australian Research Council; Canada – Natural Sciences and Engineering Research Council of Canada, Calcul Québec, Compute Ontario, Canada Foundation for Innovation, WestGrid, Digital Research Alliance of Canada, and Arthur B. McDonald Canadian Astroparticle Physics Research Institute; Denmark – Villum Fonden, Carlsberg Foundation, and European Commission; New Zealand – Marsden Fund; Japan – Japan Society for Promotion of Science (JSPS) and Institute for Global Prominent Research (IGPR) of Chiba University; Korea – National Research Foundation of Korea (NRF); Switzerland – Swiss National Science Foundation (SNSF).
Appendix A Limitations of the parametrization
The physical interpretation of Eq. 3 requires $y=[0,1]$, which means that $0<$$\langle y\rangle$ $<1$, and $\frac{dp}{dy}$ must be positive, which in turn requires that $\lambda>0$ and $\epsilon>-1$. The allowed physical space as well as these limits are shown in Fig. 14.
A special case arises when $\lambda<1$, because in that case $\lim_{y\to 0}\frac{dp}{dy}=\infty$. Because of the finite number of events in our data, and also because of presence of kinematic constraints on events with very low inelasticity, the fit might not be able to detect a sharp rise in the distribution near $y=0$. Since there is no data in this region, the integral of the distribution diverges from $1$, thus violating the probability distribution function normalization requirement. This has two undesirable consequences in our analysis. First, this means that the $\langle y\rangle$ variable in Eq. (3) loses it physical interpretation as the mean inelasticity. The second outcome is that breaking of the normalization requirement can cause the minimizer to use $\frac{dp}{dy}$ parameters to modify normalization of the data, which will absorb part of the systematic uncertainties in the study, affecting the correctness of the measurement.
To avert this problem, we adjusted the simulation to data in $\log_{10}\lambda$ instead of $\lambda$ and put a weak half-Gaussian prior on negative $\log_{10}\lambda$ with $\sigma(\log_{10}\lambda)=0.4$ on the negative side. Then, if the data prefers $\lambda<1$, the fit can still allow it with minimal bias, but if this is not the case, the prior will prevent the minimizer from getting stuck in local minima, which is especially relevant when performing a scan to determine confidence regions. The correct behavior of the fit in these cases was confirmed with Monte-Carlo tests.
Figure 14: Physically allowed and forbidden regions in the physics parameter space.
Appendix B Parameter impact on analysis sample
Figure 15: Example $\frac{dp}{dy}$ distributions for different combinations of parameters $\langle y\rangle$ and $\lambda$.
Figure 16: Modifications to the nominal expectation introduced when physics parameters in all energy bins are varied.
In our parametrization, the GENIE+HKKMS’15 model is best described by $\langle y\rangle$ $=0.43\pm 0.02$ and $\log_{10}\lambda=0.00^{+0.11}_{-0.08}$ in the energy range of interest. In order to illustrate the impact of assuming different values of $\langle y\rangle$ and $\lambda$, we calculate the expected change in number of events in the analysis binning of reconstructed energy and inelasticity. Figure 15 shows three example $\frac{dp}{dy}$ distributions with different values of $\langle y\rangle$ and $\lambda$. The comparison between the event counts corresponding to all three distributions and the expectation from GENIE+HKKMS’15 is shown on Figure 16. Figure 17 shows changes in the number of events due to modifications in systematic parameter values, while the GENIE+HKKMS’15 model is assumed for the inelasticity distribution.
Figure 17: Fractional change in the number of events compared to the nominal expectation, introduced when each nuisance parameter was increased by $1\sigma$.
In order to quantify correlations between physics and nuisance parameters in the fit, we performed a pseudo-trial test. We generated an ensemble of 400 pseudo-experiments assuming nominal values of fit parameters, but including Poisson fluctuations in each analysis bin. For each of the pseudo-experiments an inelasticity distribution fit was performed and resulting posterior parameter distributions were used to calculate Pearson correlation coefficients for each pair of fit parameters. Figure 18 shows the Pearson correlation matrix based on all 400 trials in the test.
Figure 18: Parameter correlations based on ensemble test at the best fit point.
After the fit to data was done, we also conducted a two-dimensional likelihood scan for all 15 unique pairs of physics parameters. Figure 19 shows two-dimensional likelihood scans for each pair of parameters, featuring best fit point and confidence level contours.
Figure 19: Two-dimensional likelihood scans for all unique combination of physics parameters. Black cross corresponds to the best fit point and the contours show 1, 2 and 3 $\sigma$ confidence intervals.